# Perfect supercurrent diode efficiency in chiral nanotube-based Josephson junctions

### v1

@@ Abstract
The diode efficiency takes a particularly simple form: $\eta=I_{0}/\tilde{I}_{c},$ where $\tilde{I}_{c}$ is the maximum of $\tilde{I}_{s}(\phi)$. This can be solved analytically:
$\eta$ $=\frac{2\hat{\Phi}/(LR)}{\sqrt{(2\hat{\Phi}/(LR))^{2}+\gamma^{2}}},$ (12)
@@
Here we assumed $\gamma^{-1}\neq 0$; otherwise, the diode effect vanishes ($\eta=0$). A representative calculation of $\eta$ as a function of the chiral angle $\theta$ is shown in Fig. 2(a) showing $\eta$ is suppressed at chiral angles $\theta=n\pi/2$ for $n\in\mathbb{Z}$, consistent with Ref. 35. We also observe suppression of $\eta$ at some chiral angles $\theta_{0}$ (e.g. $\sim 0.18\pi$) where $\gamma^{-1}=0$ for $\cos 2\theta_{0}=\frac{m_{0}(\kappa_{1}-\lambda)}{2\lambda m_{1}}$.
@@
We observe that $\eta$ is independent of temperature. This is due to the two quantities $(I_{c+}+I_{c-})$ and $(I_{c+}-I_{c-})$ having the same temperature scaling. We also see that $\mathrm{sgn}(\eta)=\mathrm{sgn}(B_{ext})$ for $10<\hat{\Phi}<$ 1. Both of these features stand in contrast to the superconducting diode effect predicted using the GL theory for a superconducting chiral nanotube 35 (i.e. in the absence of the junction) where the $\eta$ in that case is sensitive to $T$ and generally changes sign for $10<\hat{\Phi}<$ 1. In Ref. 35, the diode effect is calculated using a phenomenological pair breaking Cooper pair momentum along the nanotube and results in a small $\eta$ ($<0.03$). The result in Eq. (12) ignores a depairing momentum in the superconducting leads. This pair breaking mechanism is expected to play some role in actual experiments, suggesting the diode effect may have some weak $T$-dependence in reality.
@@
Evaluating the upper bounds on $\eta$ we see $\eta\rightarrow 1$ as $I_{c}/I_{0}=\gamma LR/2\hat{\Phi}\rightarrow 0,$ see Fig. 2(b). This limit is, in principle, achieved by optimizing the extrinsic contribution $\hat{\Phi}/LR$ to dominate over the intrinsic contribution $\gamma$ determined by the superfluid stiffness of the chiral nanotube (i.e. $m_{1},m_{2},\kappa_{1},\lambda$). Here $p_{y}^{(0)}$ cannot be arbitrarily enhanced by decreasing $R$ since a larger kinetic energy of the condensate eventually leads to a suppression of the superconducting state 35. Then the key parameter for optimizing $\eta$ is a small junction length $L$, as shown in Fig. 2(c-d). Figure 2(c) presents $\eta$ versus $\theta$ for a moderate junction length $L/R=10$. Here $\eta$ achieves maximum values gradually approaching $0.7$ when $\hat{\Phi}$ is increased to $0.5$. However, when $L/R=1$, Fig. 2(d) shows that $\eta$ quickly approaches nearly perfect diode efficiency ($\eta=1$) as $\hat{\Phi}$ increases beyond $0.2$. We quantify the dominance of this extrinsic contribution in terms of a geometric quantity $\eta=\sin\Theta$ where $\Theta=\arctan(I_{0}/I_{c})$, which implies $\eta$ is bounded to values $|\eta|\leq 1$. Thus, in principle it is possible to achieve perfect diode efficiency without non-equilibrium effects 33, 42, 32. Practically speaking, $\eta$ can only approach the ideal limit since $\frac{m_{1}}{m_{0}}+\cos 2\theta>0$ in this chiral nanotube system since we assumed $m_{0}<m_{1}$ (as is typically the case), but there is not a fundamental restriction against $m_{0}=m_{1}$.
@@
In this work, we presented a GL theory for a ChNt-JJ. We derived a purely orbital anomalous phase that develops across the junction when a magnetic field is applied parallel to the tube. The diode effect is suspected to be intimately connected with the anomalous phase 43, but here we find it to be independent of the SDE. We have also shown the origin of the diode effect here is a non-reciprocal persistent current that is protected by fluxoid quantization and can lead to a nearly perfect diode efficiency. While our analysis is not done for a specific material system, the general arguments are applicable to a wide variety of ChNt-JJs. An ideal setting to test SDE in chiral nanotubes is with single-walled nanotubes since effects related to vortices would be suppressed. It has been demonstrated that a thin flake of NbSe2 can induce superconductivity in carbon nanotubes and enable superconducting effects to be probed at high fields 44. This could be a useful platform for future studies of supercurrent diode effects in chiral nanotubes.
@@ Acknowledgements
J.J.C thanks Catalin D. Spataru, Wei Pan and Enrico Rossi for fruitful discussions. The work at Sandia is supported by a LDRD project. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.
Appendix A Inclusion of pair co-tunneling
In our analysis of the chiral nanotube so far we have ignored the contribution of Cooper pair co-tunneling. We will now show the combination of an anomalous phase and pair co-tunneling is not a sufficient condition for the JDE in a ChNt-JJ. To incorporate co-tunneling effects, we assume an additional contribution to the free energy in the form
[…]
$f_{kin}(\psi)$ $=\frac{1}{2m_{0}}|\mathbf{p}_{0}\psi|^{2}+\frac{1}{2m_{1}}\left(|p_{x0}\psi|^{2}-|p_{y0}\psi|^{2}\right),$ (14)
where $f_{kin}$ is the kinetic part of pair condensate free energy density and for $\Gamma\in\mathbb{R}$ proportional to the pair co-tunneling probability across the JJ. Then the supercurrent density of a short ballistic chiral nanotube-based JJ is given by
$\mathbf{J}$ $=\frac{2e\hbar}{i}\left(1+2\Gamma f_{kin}(\psi)\right)\begin{pmatrix}\rho_{1}&&\rho_{3}\\ \rho_{3}&&\rho_{2}\end{pmatrix}\mathbf{j}.$ (15)
Importantly, we see the form of Eq. (13) creates the same superfluid stiffness anisotropy in the pair and pair co-tunneling channels. The CPR of the small-diameter tube in this case is
$I_{s}(\phi)=$ $\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}\sin\left(\phi+\phi_{0}\right)$
$\cdot\left[1+4\hbar^{4}\Gamma\frac{\psi_{\infty}^{2}\rho_{1}}{L^{2}}\left(1-\cos\left(\phi+\phi_{0}\right)\right)\right]$ (16)
$=$ $~{}I_{c}\left(1+2\delta_{\Gamma}\right)\sin\tilde{\phi}-\delta_{\Gamma}\sin 2\tilde{\phi},$ (17)
where $\delta_{\Gamma}=2\hbar^{4}\Gamma\psi_{\infty}^{2}\rho_{1}/L^{2}$. Clearly, there is no diode effect despite an anomalous phase and pair co-tunneling. This is because by propagating the same anisotropy in the pair current channel to the pair co-tunneling channel, the same anomalous phase emerges in both channels suppressing any interference effect between the two channels. If an isotropic pair co-tunneling term had been includes instead of $f_{kin}$, then there may be some interference between the channels resulting in a JDE but the GL equations requires a numerical solution.

### v2

@@ III.2 Small nanotube diameter limit
Now let’s characterize the diode effect using Eq. (10). The diode efficiency takes a particularly simple form: $\eta=I_{0}/\tilde{I}_{c},$ where $\tilde{I}_{c}$ is the maximum of $\tilde{I}_{s}(\phi)$. This can be solved analytically:
$\eta$ $=\mathrm{sgn}(\gamma)\frac{2\hat{\Phi}/(LR)}{\sqrt{(2\hat{\Phi}/(LR))^{2}+\gamma^{2}}},$ (13)
@@
Here we assumed $\gamma^{-1}\neq 0$; otherwise, the diode effect vanishes ($\eta=0$). A representative calculation of $\eta$ as a function of the chiral angle $\theta$ is shown in Fig. 4(a) showing $\eta$ is suppressed at chiral angles $\theta=n\pi/2$ for $n\in\mathbb{Z}$, consistent with our numerical calculations and Ref. 9. We also observe suppression of $\eta$ at some chiral angles $\theta_{0}$ (e.g. $\sim 0.18\pi$) where $\gamma^{-1}=0$ for $\cos 2\theta_{0}=\frac{m_{0}(\kappa_{1}-\lambda)}{2\lambda m_{1}}$.
@@
We observe that $\eta$ is independent of temperature. This is due to the two quantities $(I_{c+}+I_{c-})$ and $(I_{c+}-I_{c-})$ having the same temperature scaling. This stands in contrast to the superconducting diode effect predicted using the GL theory for a superconducting chiral nanotube 9 (i.e. in the absence of the junction) where the $\eta$ in that case is sensitive to $T$. In Ref. 9, the diode effect is calculated using a phenomenological pair breaking Cooper pair momentum along the nanotube and results in a small $\eta$ (<0.03<0.03). The result in Eq. (13) ignores a depairing momentum in the superconducting leads. This pair breaking mechanism is expected to play some role in actual experiments, suggesting the diode effect may have some weak $T$-dependence in reality.
IV Discussion
@@
Evaluating the upper bounds on $\eta$ we see $\eta\rightarrow 1$ as $I_{c}/I_{0}=\gamma LR/2\hat{\Phi}\rightarrow 0,$ see Fig. 4(b). This limit is, in principle, achieved by optimizing the extrinsic contribution $\hat{\Phi}/LR$ to dominate over the intrinsic contribution $\gamma$ determined by the superfluid stiffness of the chiral nanotube (i.e. $m_{1},m_{2},\kappa_{1},\lambda$). Here $p_{y}^{(0)}$ cannot be arbitrarily enhanced by decreasing $R$ since a larger kinetic energy of the condensate eventually leads to a suppression of the superconducting state 9. Then the key parameter for optimizing $\eta$ is a small junction length $L$, as shown in Fig. 4(c-d). Figure 4(c) presents $\eta$ versus $\theta$ for a moderate junction length $L/R=10$. Here $\eta$ achieves maximum values gradually approaching $0.7$ when $\hat{\Phi}$ is increased to $0.5$. However, when $L/R=1$, Fig. 4(d) shows that $\eta$ quickly approaches nearly perfect diode efficiency ($\eta=1$) as $\hat{\Phi}$ increases beyond $0.2$. We quantify the dominance of this extrinsic contribution in terms of a geometric quantity $\eta=\sin\Theta$ where $\Theta=\arctan(I_{0}/I_{c})$, which implies $\eta$ is bounded to values $|\eta|\leq 1$. Thus, in principle it is possible to achieve perfect diode efficiency without non-equilibrium effects 41, 50, 40. Practically speaking, $\eta$ can only approach the ideal limit since $\frac{m_{1}}{m_{0}}+\cos 2\theta>0$ in this chiral nanotube system since we assumed $m_{0}<m_{1}$ (as is typically the case), but there is not a fundamental restriction against $m_{0}=m_{1}$ where we also found numerically in Fig. 2(c) that $\eta$ approaches unity. The key ingredients we identify in our work to realize perfect SDE in a weak link are: (i) an anisotropic superfluid stiffness and (ii) fluxoid quantization associated with the applied axial magnetic field which gives rise to a non-reciprocal persistent current across the JJ. The existence of a non-reciprocal persistent current in the JJ is not constrained to the vicinity of the superconducting critical temperature; thus, we expect our analysis to also be relevant at low temperatures.
V Conclusion
In this work, we presented a GL theory for a ChNt-WL. We derived a purely orbital anomalous phase that develops across the junction when a magnetic field is applied parallel to the tube. The diode effect is suspected to be intimately connected with the anomalous phase 51, but here we find it to be independent of the SDE. We have also shown the origin of the diode effect here is a non-reciprocal persistent current that is protected by fluxoid quantization and can lead to a nearly perfect diode efficiency. We found this non-reciprocal persistent current can lead to perfect SDE in a weak link—a counter-intuitive result in the context of the basic theory of the dc Josephson effect where persistent currents cannot contribute to the Josephson energy. In our analysis, we neglected the effect of thermally excited quasiparticles because the key ingredients we identify in our work to realize perfect SDE in a JJ are unchanged: (i) an anisotropic superfluid stiffness and (ii) fluxoid quantization associated with the applied axial magnetic field which gives rise to a non-reciprocal persistent current across the JJ. The presence of thermally excited Bogoliubov quasiparticles will certainly affect the voltage-current characteristic of the weak link, but a central result of our work is that nearly perfect SDE can be achieved if the extrinsic non-reciprocal persistent current dominates over the intrinsic Josephson current in the junction. The phase-independent persistent current arises specifically in a geometry where fluxoid quantization is a constraint on the gauge-invariant superconducting phase and a combination of mirror and time-reversal symmetries are broken, such as in asymmetric SQUIDs with a finite self-inductance.
We want to emphasize that the key finding in our work is a resolution to the problem of achieving perfect diode efficiency in a Josephson system. In our work, we overcome this challenge by showing a persistent current is found in the chiral nanotube weak link. Despite starting with the same free energy functional as Ref. 9, our calculation is distinct from that of Ref. 9 where they calculate extrema of the supercurrent density with respect to variations in a phenomenological depairing momentum. In our approach, we calculate extrema of the CPR, which does not consider a depairing momentum. The superconducting diode efficiency reported in Ref. 9 has a maximum value less than 0.03. Using our model for a weak link with comparable parameters, we find a nearly perfect diode efficiency (>>0.99). This is the first step in developing a theoretical understanding of a perfect dc SDE in a chiral nanotube-based Josephson junction, paving the way for future experiments in this direction.
The mechanism for approaching perfect SDE is also distinguished from prior work in the literature on spin-orbit coupling-based mechanisms. At the level of Ginzburg-Landau theory, spin-orbit coupling-based mechanisms arise from a Lifshitz invariant in the free energy density to account for broken spatial inversion symmetry 45, 49. In contrast, in our work the free energy density is invariant under spatial inversion. Secondly, while both chiral nanotube and Rashba JJs have a broken mirror symmetry across the junction, the SDE in Rashba systems is usually described in terms of the Zeeman effect while the SDE in the chiral nanotube is associated with an orbital magnetic effect. A consequence of gauge invariance in a nanotube geometry is fluxoid quantization (e.g. Little-Parks effect) which is crucial for nearly perfect diode efficiency in our work. A theoretical proposal to use a persistent current to achieve perfect diode efficiency in a proximity-induced superconducting material with band asymmetry was reported previously 52. We show a non-reciprocal persistent current can develop in a weak link and lead to perfect SDE, opening the door to further study of Josephson devices where unusual persistent currents dictate the non-reciprocal transport properties.
@@ V Conclusion
While our analysis is not done for a specific material system, the general arguments are applicable to a wide variety of ChNt-WLs. An ideal setting to test SDE in chiral nanotubes is with single-walled carbon nanotubes. Long coherence length can be achieved with common conventional superconductors like Al ($\sim$1 um). Devices with relatively low trapped charge disorder can be made by suspending carbon nanotubes across a trench 53 or by encapsulating tubes with a hexagonal boron nitride thin film 54, similar to graphene-based devices. In the case of a single-walled carbon nanotube, an axial magnetic field has minimal magnetic flux penetrating the tube so that effects due to vortices are minimal. It has been demonstrated that a thin flake of NbSe2 can induce superconductivity in carbon nanotubes and enable superconducting effects to be probed at high fields 54. This could be a useful platform for future studies of supercurrent diode effects in chiral nanotubes.
@@ Acknowledgements
J.J.C thanks Catalin D. Spataru, Wei Pan and Enrico Rossi for fruitful discussions. The work at Sandia is supported by a LDRD project and the U.S. Department of Energy, Office of Science, under the Accelerate Innovations in Emerging Technologies Program. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.